OpenStudy (anonymous):

MEDAL AND FAN: PLEASE HELP ME!!! What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (anonymous):

is this like graph wise?

OpenStudy (anonymous):

or elimination

OpenStudy (photon336):

you could also eliminate one of the variables, so you can combine the two equations and solve for one variable.

OpenStudy (anonymous):

x-y=-3 x+3y=5 So add -3+5 then eliminate the x's u have left: -y and 3y with 2 sp add them up 4y=2 and u got the rest

OpenStudy (anonymous):

and thats the x part

OpenStudy (photon336):

well you could multiply the entire top equation by -1 and then add the top and bottom equations. for instance you would have -x+x = 0. meaning x would cancel ou.t